OpenStudy (anonymous):

Please help! Will fan and medal! Use substitution for this problem. -3x+y=-3 y=x-3

Mehek (mehek14):

substitute y = x - for y in the first equation

Mehek (mehek14):

y = x - 3*

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i think i understand it now is the answer (2,3)

Mehek (mehek14):

no..

OpenStudy (anonymous):

oh

Mehek (mehek14):

it should be (0,-3) when you substitute you solve for x then you put the value of x into the equation and get y

OpenStudy (anonymous):

No disrespect but I felt pretty confident with my answer :(

Mehek (mehek14):

-3x + x - 3 = -3 -2x = 0 x = 0 use the second equation y = x - 3 put 0 for x y = 0 - 3 y = -3 (0,-3)

OpenStudy (anonymous):

ok